OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, –1) and (–2, –5)?

OpenStudy (the_fizicx99):

M = y2-y1/x2-x1

OpenStudy (mathmale):

Haruk: Please type out the formula for slope of a straight line and then use it to find the slope of the given line through those two points.

OpenStudy (luigi0210):

I already answered this in chat for you dang it ლ(ಠ益ಠლ) But use the formula: \[\LARGE m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] Now finish it: \[\LARGE m=\frac{-5+1}{-2-3}\]

OpenStudy (anonymous):

-4/-5 ??????????? @Luigi0210

OpenStudy (the_fizicx99):

Two negatives!!

OpenStudy (the_fizicx99):

as my uncoordinated friend said, ლ(ಠ益ಠლ)

OpenStudy (luigi0210):

Just 4/5

OpenStudy (the_fizicx99):

ლ(ಠ益ಠლ)